Moon Software Inc. is planning to issue two types of 25-year, noncallable bonds to raise a total of $6 million, $3 million from each type of bond. First, 3,000 bonds with a 10% semiannual coupon will be sold at their $1,000 par value to raise $3,000,000. These are called "par" bonds. Second, Original Issue Discount (OID) bonds, also with a 25-year maturity and a $1,000 par value, will be sold, but these bonds will have a semiannual coupon of only 6.75%. The OID bonds must be offered at below par in order to provide investors with the same effective yield as the par bonds. How many OID bonds must the firm issue to raise $3,000,000? Disregard flotation costs, and round your final answer up to a whole number of bonds.

a. 4,266
b. 3,370
c. 3,285
d. 4,906
e. 4,479
Answer #1

Price of bond is equal to present value of all future coupon payments and the principal amount

= 1000*6.75%*1/2*PVAF(5%, 50 periods) + 1000*PVF(5%, 50 periods)

= 33.75*18.2559 +1000*0.0872

= $703.34

Number of bons required to be issued = 3,000,000/703.34

= 4,265.36 bonds

i.e. 4,266 bonds

i.e. a